3. During the execution of a reversible non-flow process the work is -156.2 KJ if V 1 =0.845 and the

pressure varies as P=-730 V +690 KPa, where volume V is in m 3 . Find the final volume.


1
Expert's answer
2021-09-15T16:16:43-0400

"Work = -156.2 \;kJ = -43.34 \;Wh \\
V_1 = 0.845 \\
P = -730V + 690 \;kPa \\
W = \int VdP = V(P_1 -P_2) \\
P_1 = -730 \times 0.845 + 690 = 73.15 \;kPa \\
P_2 = -730V_2 + 690 \\
W=V(P_1-P_2) \\
-43.34 = 0.845(73.15 -P_2) \\
P_2 = 124.439 \;kPa \\
124.439 = -730V_2 + 690 \\
V_2 = 0.775 \;m^3"